OpenStudy (aravindg):

find the eqn of normal to the curve y=(1+x)^y +sin inverse (sin^2 x) at x=0

OpenStudy (aravindg):

@apoorvk , @Mani_Jha

OpenStudy (aravindg):

@asnaseer , @satellite73 , @TuringTest

OpenStudy (mani_jha):

Find dy/dx at x=0 Let the point be (0,b) Slope of normal: \[-1/(dy/dx)\] Equation of normal: \[(y-b)=(-1/(dy/dx))(x-0)\] Got it?

OpenStudy (aravindg):

hmm. w8

OpenStudy (experimentx):

this doesn't look easy.

OpenStudy (asnaseer):

Mani_jha is correct - spot on!

OpenStudy (aravindg):

i will try tht and let u knw if i get stuck..anyway thx all

OpenStudy (experimentx):

I think the question is more like, solve for y, then find dy/dx put the values of x,y on dy/dx then calculate the normal.

OpenStudy (experimentx):

looks like it was that point mani mentioned was 0,1 I got crazy value for dy/dx ... please check it out.

OpenStudy (asnaseer):

experimentX: you can just use implicit derivative - no need to solve for y first.

OpenStudy (experimentx):

still, the dy/dx will be interms of y, so y is needed at that point,

OpenStudy (asnaseer):

dy/dx will most likely be near 0 at x=0 by the looks of the curve

OpenStudy (experimentx):

yeah ... but i got 1, y=1, ... looks like there must some error.

OpenStudy (experimentx):

*m=1

OpenStudy (asnaseer):

\[y=(1+x)^y+\sin^{-1}(\sin^2(x))\]\[y'=y(1+x)^{y-1}*y'*1+\frac{\sin(2x)}{\sqrt{1-\sin^4(x)}}\]which leads to:\[y'=\frac{\sin(2x)}{(1+y(1+x)^{y-1})\sqrt{1-\sin^4(x)}}\]

OpenStudy (asnaseer):

unless I have made a mistake somewhere, this gives y'=0 at x=0

OpenStudy (experimentx):

yeah ... it's zero, looks like i made mistake in taking dy/dx

OpenStudy (experimentx):

does (x+1)^y work with chain rule??

OpenStudy (asnaseer):

I believe so

OpenStudy (aravindg):

ready for the next qn?

OpenStudy (experimentx):

I thought it was some what like x^x, or (1+x)^f(x) <--- though I am not quite sure on this.

OpenStudy (asnaseer):

just checked with the wolf - you could be right experimentX: http://www.wolframalpha.com/input/?i=differentiate+y%3D%281%2Bx%29^y+with+respect+to+x

OpenStudy (experimentx):

i hope so.. still not getting the value of m. let me revise.

OpenStudy (asnaseer):

the derivative at x=0 works out to be equal to y.

OpenStudy (experimentx):

?? I got the same result. dy/dx = 1 + 0

OpenStudy (experimentx):

might be due to this crazy function http://www.wolframalpha.com/input/?i=graph+y+%3D+%281%2Bx%29%5Ey